OpenStudy (anonymous):

Can someone explain to me the rules of adding and subtracting negative numbers

OpenStudy (ledah):

Positive + Positive = Positive: 5 + 4 = 9 Negative + Negative = Negative: (- 7) + (- 2) = - 9 Sum of a negative and a positive number: Use the sign of the larger number and subtract (- 7) + 4 = -3 6 + (-9) = - 3 (- 3) + 7 = 4 5 + ( -3) = 2 Subtracting Rules: Negative - Positive = Negative: (- 5) - 3 = -5 + (-3) = -8 Positive - Negative = Positive + Positive = Positive: 5 - (-3) = 5 + 3 = 8 Negative - Negative = Negative + Positive = Use the sign of the larger number and subtract (Change double negatives to a positive) (-5) - (-3) = ( -5) + 3 = -2 (-3) - ( -5) = (-3) + 5 = 2 Hope this helps

Parth (parthkohli):

\[\Large \color{MidnightBlue}{\Rightarrow -x - (-y) = -x + y }\] \[\Large \color{MidnightBlue}{\Rightarrow -x + (-y) = -x - y }\] Goes like this.

Parth (parthkohli):

You asked for only negative numbers, so here it is.

Parth (parthkohli):

Also, \[\Large \color{MidnightBlue}{\Rightarrow x - y = -(y - x) }\]

OpenStudy (anonymous):

if you add two negatives together, you get a smaller number positive+negitive =positive-positive negitive -negitive=positive - negitive positive - negitive=positive +positive negitive -positive=negitive+negitive

OpenStudy (anonymous):

I gave you both

OpenStudy (ledah):

I keep going overboard on my explanations